OpenStudy (anonymous):

how much heat is given off by 1000 kg of iron when it solidifies at its melting point? A. 2.66 * 10^2 J B. 3.76 * 10^-3 J C. 2.66 * 10^7 J D. 2.66 * 10^8 J ** i will draw a table!

OpenStudy (anonymous):



OpenStudy (yadu123):

how many questions is this reveiw packet

OpenStudy (anonymous):

i got access to the tools again!! :P

OpenStudy (michele_laino):

here we have to compute this: (heat of fusion of iron) times (mass of iron)

OpenStudy (anonymous):

okay so 1000 times _____ ? not sure what to multiply it by :/

OpenStudy (yadu123):

time the heat of fusion for iron

OpenStudy (michele_laino):

\[{10^3} \times 2.66 \times {10^5} = ...Joules\]

OpenStudy (anonymous):

oh okie! so we get this? 266000000? so our solution is choice D?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

ooh yay! thank you!!:D

OpenStudy (michele_laino):

:D